OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -12 and 768, respectively.

OpenStudy (anonymous):

an = 3 • (-4)^n + 1 an = 3 • 4^n - 1 an = 3 • (-4)^n - 1 an = 3 • 4^n

OpenStudy (anonymous):

help anyone?????